Bremerton Auto Wrecking - Gorst Creek Site

Notices

All Notices [5] rss feed
Categories
Notices  Posted Category
Removal Complete - Restoration Progressing 9/27/2017 Site Update

The last of the landfill waste was excavated on July 19, 2017 and on July 26, 2017 the last of the landfill waste was shipped off site for treatment or disposal. A total of 337,534 tons of waste was removed.

Stream channel construction is complete, the slope soils have been amended with top soil and mulch. Workers are currently planting native vegetation and placing woody debris on the slopes. Major site work is expected to be complete in October 2017 with additional planting, monitoring and repairs occurring as needed this coming winter.

September 2017

Picture taken from the same location as pictures in previous notices (on the east side of the landfill looking northwest, towards the downstream end of the landfill).



Update April 15 - June 20, 2017
Deepest Part of Landfill Removed
6/23/2017 Site Update

Excavation activities on the landfill have been continuing since early April.  This has included the removal of the temporary conveyance channel which was in place over the Winter to allow the creek flow over the remaining landfill material/

In May 2017, excavators began removing the culvert which was under the landfill.  So far about one third of the culvert has been removed from the downstream end of the landfill. This means that the part of the landfill where the debris was the deepest has been removed - debris and hazardous substances were roughly 60 feet deep in this area.

About sixty percent of the landfill surface has had all of the visible debris removed and has been sampled to confirm that soil and sediment are not hazardous. All landfill debris must be removed and all sediment must be sampled to confirm it is clean before restoration can begin.

So far 275,000 tons of material have been removed; 6,800 tons were recycled and 2,000 tons were hazardous.

The photos below were all taken on the east side of the landfill looking northwest, towards the downstream end of the landfill.

June 2016

Approximately one month into excavation.

September, 2016

Excavators will soon construct an overflow channel to allow water flow over the landfill during the winter.

June, 2017

The valley is taking shape. Workers are pumping stormwater from the top of the landfill. Once it is dry the last of the landfill debris can be removed



Update Mar 27-April 15, 2017. Remobilization: Excavation and Off Site Disposal Resume. 4/18/2017 Site Update
  • EPA, USCG, START, and ERRS personnel remobilized to the site on Monday, March 27. ERRS mobilized a partial crew, with the remainder of their staff scheduled to arrive over the next couple of weeks.

  • Staff performed a general health and safety review.

  • Set up tasks included mobilizing equipment, Generators, instrument field deployment procedures, Set up air stations, and equipment maintenance and tracking.  

  • Re-grade and maintain on site and transportation haul roads.

  • Excavation of landfill and filling of stockpiles resumes on Mar 29 

  • Sampling of first full stockpile was on Mar 31.

 Landfill Material Hauling to Stockpile and Off Site Disposal/Recycling totals will be provided under documents with tables/spreadsheets updated with and monthly totals.



Winter Ops SW Weekly Inspections Nov 2016 - Mar 2017 4/3/2017 Site Update

 
 

● SWPPP inspection was performed at least weekly through the Winter.  And also following major rainfall events.

This included inspection of all the on site BMPs in place and inspection of the on site stormwater pond.

Also inspected weekly was the upstream impoundment, the conveyance channel and lower creek outfall.     

· ● During the week of October 27 and 28, ERRS was on site to perform slope and drainage maintenance including installation of coir fabric and straw wattles, placement of mulch (hog fuel), and reinforcement of silt fences. These features were inspected and no deficiencies were observed.

Samples for turbidity levels in the lower portion of Gorst creek were also collected and analyzed on site each week. 

     ●  Week of Oct 17-Grass coverage in the landfill continued to improve with blades about 1.5 inches tall. Rills were stable and did not appear to be extending. The repairs made in late October were in good condition.


               
  • During the week of Nov 28 ERRS was on site to perform slope and drainage maintenance. All repairs were inspected and are in good condition The North slope areas where rills and channels had formed were raked out and covered with coir fabric and straw wattles. Areas sloping to the site exterior where mulch (hog fuel) was not placed have logs and stumps to hold soil in place. Other areas are sloped toward the interior of the site to prevent runoff.

 

Tuesday, January 24              

  • Planting is scheduled in the test plot area on the north slopes near the AAW pond. The AAW pond is ¾ full but not discharging. The stockpile ponds are in good condition with no discharging observed.

 

Tuesday, January 24- Friday, January 27: Test Plot Planting

  • 1,100 bare-root and potted plants were placed in the 0.6 acre test plot during a four-day period in late January
    • Included 400 trees comprised of 4 different species
    • Included 700 shrubs comprised of 7 different species
  • The test plot was divided into three sections (T1, T2, and T3)
    • T1 included compost only
    • T2 included compost and some topsoil amendment where compost was less than 3 inches in depth
    • T3 included compost and hog fuel (mulch)
  • Slow release 14-14-14 fertilizer granules were added to each hole
  • Small depression was formed around each plant for water catchment
  • The test plot feature was incorporated into the weekly storm water inspection to assess for impact from wildlife and precipitation (or lack thereof).

 




Site Update Sept. 26 - Oct 29, 2016 Conveyance Channel Construction - STORM Plan Implementation 3/28/2017 Site Update

 

Monday, September 26 - 27

● Excavation from landfill and filling of stockpiles continued.

Continue  to line the downstream creek outfall area with concrete pieces and PVC liner in preparation for the winter STORM operations. ERRS also began to excavate the STORM surface water channel through the center of the remaining landfill area.
● ERRS removed the pumps, generators, and piping that had been used for the creek surface water bypass system.
● ERRS began to move the chain-link fence between the landfill and Highway 3 to the perimeter of the excavation area in preparation of winter STORM operations.

● START submitted the updated STORM plan with 2016-2017 winter planting recommendations to EPA and ERRS.
● START and USCG performed air monitoring with AreaRAEs, MultiRAEs, DataRAMs, and DustTraks and collected air samples for asbestos and lead. Air monitoring results were below action levels and within normal limits.
● Off-site laboratory results were received for lead air samples collected previously; all results were below site action levels.

 

● EPA ERT staff  arrived on site to assess the seven cylinders.   START mobilized the EPA Region 10 Level A truck to support the cylinder assessment and neutralization, and the USCG provided health and safety oversight. Six cylinders were related to industrial firefighting and confirmed to be empty; the valves on these cylinders were disabled and the cylinders were set aside for recycling. The remaining cylinder of suspect acetylene was prepared for processing on the following day.

● Oct 4 During the afternoon operations meeting, the following decisions were made: PPE requirements were downgraded to modified Level D unless task-specific job hazards required upgrading; no additional fixed or personal air sampling will be performed during the remainder of the current field event; and, no electronic personal dosimeters (EPDs) will be issued during the remainder of the current field season.

 

● ERT staff prepared to remotely drill the cylinder of suspect acetylene. The cylinder was successfully drilled. A de minimis volume of residual acetylene was present in the cylinder and allowed to vent to atmosphere. Due to concern that the cylinder may contain asbestos, it was set aside to be shipped off site for disposal as asbestos-containing material (ACM). START and USCG provided support to ERT including air monitoring, site setup, and tear down activities.

● START shipped five cases of equipment used during the cylinder operations to the ERT warehouse in Las Vegas, NV. 

 

● During the afternoon operations meeting, the following decisions were made: AreaRAEs will no longer be deployed; the remaining Air Stations 1, 3, and 5 will be decommissioned; no particulate monitors will be deployed for the remainder of the field season, and therefore all Viper-related activities will be discontinued.